In a Stirling cycle, the pressure and temperature at the start of isothermal compression are 1 bar and 40°C. The compression ratio is 7 and the maximum temperature in the cycle is 1200°C. Determine;

i) The maximum pressure and the mean effective pressure

ii) The thermal efficiency of the cycle with and without a regenerator



1
Expert's answer
2021-07-19T15:43:30-0400

Gives

"P_1=1bar=10^5pa\\T_1=40°c\\T_2=1200°c\\r=\frac{V_1}{V_2}=7"

We know that gas eqution

PV=nRT

"\frac{P_1V_1}{P_2V_2}=\frac{T_2}{T_1}"

Adibatic process

"TV^{r-1}=constant"

We can written as

"P_2=P_1(\frac{V_1}{V_2}\times\frac{T_2}{T_1})"

"P_2=10^5\times(7\times\frac{1200}{40})=210pa"

Thermal efficiency

"\eta=1-\frac{T_2}{T_1}"

"\eta=1-\frac{313}{1473}=0.78"

"\eta" %=78%

Thermal efficiency 78%